OpenStudy (anonymous):

SOLVE THE EQUATION (Y-2)(Y-6)=0 Y=

OpenStudy (anonymous):

@AkashdeepDeb

OpenStudy (anonymous):

y = 2, y = 6

OpenStudy (mathstudent55):

The zero-product property states: if a product equals zero, then at least one of the factors is zero. y -2 and y - 6 are two factors. When multiplied together, the product equals zero. That can only happen if y - 2 = 0 or if y - 6 = 0. Solve each equation and separate the answer with the word "or." y - 2 = 0 or y - 6 = 0

OpenStudy (anonymous):

@mathstudent55 i just i did a problem and it was (x-3)(3x+1)=0 and the answer was 3,-1/3 i really dont know how to do these times of problems

OpenStudy (mathstudent55):

This is similar to the problem here. Any time you have a product of factors equal to zero, set each factor equal to zero. For the original problem of this post, this is how the solution should look: (y - 2)(y - 6) = 0 y - 2 = 0 or y - 6 = 0 y = 2 or y = 6

OpenStudy (mathstudent55):

For the second problem you just posted, you solve it similarly: \((x-3)(3x+1)=0\) \(x - 3 = 0\) or \(3x + 1 = 0\) \(x = 3 \) or \(3x = -1\) \(x = 3 \) or \(x = -\dfrac{1}{3} \)